OpenStudy (kj4uts):

1. When calcium carbonate is heated vigorously, calcium oxide and carbon dioxide are formed. CaCO3(s) → CaO(s) + CO2(g) When 10.0 g of calcium carbonate is heated, 4.00 g of carbon dioxide is obtained. Calculate the percent yield. A. 90.9 B. 40.0 C. 14.0 D. 400 2. Upon heating, ammonium nitrate decomposes as shown below. NH4NO3(s) → N2O(g) + 2 H2O(g) If a sample of ammonium nitrate yields 0.50 mole H2O, how many moles of N2O are formed? A. 0.50 mole B. 0.125 mole C. 1.0 mole D. 0.25 mole Please explain. Thank you!

OpenStudy (kj4uts):

@SapphireMoon I was wondering if you might be able to help me with these problems.

OpenStudy (kj4uts):

For the first one is it like 4/10=0.4*100=40?

OpenStudy (kj4uts):

For the second one I was thinking if you have 2 H2O(g) and it's 0.50 mole and you have N2O I was thinking it is half so 0.25 because the H2O has the 2 (I think thats the coefficient)?

OpenStudy (sapphiremoon):

No, it's not quite that simple, I'm afraid. We'll have to do mass to mole to mass conversion. We have 10 g of calcium carbonate, which has a molar mass of 100g/mol, and 1 mol of calcium carbonate makes 1 mol of carbon dioxide, which has a molar mass of 44.01g/mol. We can put this in a lovely factor-label equation: \[10.0 gCaCO_{3} \times \frac{ 1 mol }{ 100 g } \times \frac{ 1 mol CO_{2} }{ 1 mol CaCO_{3} } \times \frac{ 44.01 g CO_{2} }{ 1 mol } = 4.40g CO_{2}\] Which gives us a percent yield of 90.9%. For the second problem, you're entirely correct. If 1 mol ammonia yields 1 mol nitrous? oxide and 2 mol water, than it has to be 0.25 mol ammonia that yields 0.25 mol nitrous? oxide and 0.5 mol water.

OpenStudy (kj4uts):

Wow, I was way off on the first one, thank you very much for taking the time to write all that out it helps me to see it visually.

OpenStudy (sapphiremoon):

You're quite welcome! My chem teachers showed us that way so we'd see what was going on and never lose our units.

OpenStudy (kj4uts):

I have a midterm coming up, is the periodic table going to be my "best friend" when it comes to finding all this data?

OpenStudy (sapphiremoon):

Yeah, you're going to want to memorize the atomic symbols and numbers for at least the first three rows, my AP and Honors Chem teachers had us do four, and my AP Chem teacher had us memorize the molar masses for hydrogen, carbon, nitrogen, oxygen, fluorine, sodium, magnesium, aluminum, phosphorus, sulfur, chlorine, potassium, and calcium. Also for a couple of compounds like carbon dioxide and water of course.